### v1

# Out-of-sample gravity predictions

Abstract
Gravity equations are often used to evaluate counterfactual trade policy scenarios, such as the effect of regional trade agreements on trade flows. In this paper, we argue that the suitability of gravity equations for this purpose crucially depends on their out-of-sample predictive power. We propose a methodology that compares different versions of the gravity equation, both among themselves and with machine learning-based forecast methods such as random forests and neural networks. We find that the 3-way gravity model is difficult to beat in terms of out-of-sample average predictive performance, further justifying its place as the predominant tool for applied trade policy analysis. However, when the goal is to predict individual bilateral trade flows, the 3-way model can be outperformed by an ensemble machine learning method.
JEL Classification: C45, C53, C63, F14
Keywords: Causal inference, forecasting, gravity equation, machine learning, out-of-sample, trade costs
1 Introduction
Gravity equations are by far the most popular tool in international trade for evaluating the effects of policy interventions such as regional trade agreements (RTAs). In its most basic form, the gravity model states that exports from country $i$ to country $j$ at time $t$ are a log-linear function of exporter and importer characteristics (e.g., GDP) and bilateral variables such as distance.
The effects of RTAs and other policy interventions are usually evaluated by including an additional dummy variable for whether two countries $i$ and $j$ have an RTA in place. The coefficient on this variable then tells us by how much an agreement increases trade, over and above what is predicted by the other variables. [^1] Note that this coefficient estimate only provides a partial equilibrium effect. That is, the gravity equation is assumed to be the data generating process of the underlying counterfactual outcome in which the two countries do not have a trade agreement.
A key justification for why gravity equations are well suited for providing this counterfactual is that they have very high in-sample explanatory power. For example, even in their most basic form, gravity equations often achieve an $R^{2}$ of $70$-$80\%$ when estimated on datasets for aggregate country-level trade flows. In addition, gravity equations are commonly estimated with a large set of fixed effects to capture hard-to-observe determinants of trade flows such as local price indices, leading to an even higher $R^{2}$.
However, such statements about the empirical performance of gravity equations usually relate to in-sample predictive performance, whereas out-of-sample predictive performance has not received much attention. [^2] Recent exceptions include Ruzicska et al. (2024) and Kiyota (2025) However, none of these contributions make the connection between out-of-sample performance and causal inference, which is the key focus of our paper. Of course, good in-sample predictive performance does not necessarily imply good out-of-sample performance if the former is achieved by overfitting statistical models. [^3] Specifically, the danger of overfitting a model in-sample is that it approximates random noise in the data. Once we move to a different sample with a different realization of random noise, it will perform poorly.
In this paper, we propose a methodology to compare the predictive performance of different approaches. We then use it to carry out an evaluation of the out-of-sample forecasting performance of the gravity equation and of possible alternatives based on flexible statistical techniques from the machine learning literature, such as random forests and neural networks.
We believe that providing evidence on the relative out-of-sample predictive performance of the gravity equation is important for two reasons. [^4] Note that we will be using the terms predicting and forecasting interchangeably throughout this paper. First, having high-quality statistical models for predicting trade flows is directly relevant for policymaking. This often requires making such forecasts, for example to predict the potential trade flows for a given pair of countries. [^5] For example, the International Trade Centre has developed a so-called ‘export potential assessment methodology’ in order to help countries identify promising products for export promotion activities (see ITC, 2025). The more accurate the forecast of potential trade flows between two countries, the better such export promotion activities can be targeted.
Secondly, and more importantly, a good out-of-sample predictive performance is crucial for obtaining good estimates of the causal impact of policy interventions such as RTAs. This is because the suitability of the gravity equation as a baseline counterfactual depends on its predictive performance out-of-sample, not its performance in-sample (see Section 2 below for more details on this point). Put differently, if we can find an alternative statistical technique that provides better out-of-sample predictions for trade flows, then maybe this alternative technique should take the place of existing gravity equation techniques to generate counterfactual outcomes and evaluate the impact of policy interventions such as new RTAs. This in turn might lead to a re-examination of the reliability of a large and growing literature in areas such as the trade impact of RTAs, currency unions, and WTO membership (Baier and Bergstrand, 2007, Rose, 2001, and Rose, 2004).
To evaluate the out-of-sample predictive ability of gravity equations, we compare their performance with that of a number of alternative statistical techniques from the machine learning literature, such as random forest and neural networks. These techniques provide a useful point of comparison because they were designed with out-of-sample forecasting performance in mind. Thus, if the gravity equation does well in terms of this criterion against these alternative techniques, we can be confident that its current dominant position as an empirical tool for making forecasts and evaluating policy interventions is indeed merited.
With these goals in mind, we use a standard gravity dataset based on UN Comtrade to compare the traditional gravity equation, and more recent variants such as the 2-way and 3-way gravity models, with the following machine learning techniques: random forests, neural networks, gradient boosting, and an ensemble method based on these three techniques.
Our results suggest that the gravity equation performs well against these highly flexible forecasting tools, suggesting that it captures essential features of the data generating process of international trade flows. However, this is only true for the 3-way gravity model, which controls for exporter-time, importer-time, and bilateral fixed effects. By contrast, the traditional gravity equation (no fixed effects) and the 2-way gravity model (exporter-time, importer-time fixed effects only) can be outperformed by some of the machine learning methods we consider. This underlines the importance of bilateral fixed effects in the 3-way gravity model. Given that the 3-way model is already considered to be the state-of-the-art version of the gravity equation (see Yotov et al., 2016), we see our results as a further justification for its use.
Our research contributes to several strands in the international trade literature. First, we contribute to the literature on using gravity equations as a tool to evaluate trade policy interventions (see Goldberg and Pavcnik, 2016, and Harrison and Rodríguez-Clare, 2010, for an overview). Our results show that the gravity equation does indeed merit its place as the dominant empirical tool in this area, albeit only in its most recent incarnation as the 3-way gravity model.
Second, we contribute to the voluminous literature examining how to best estimate and interpret gravity equations (see Head and Mayer, 2014 for an overview). We contribute to this literature by pointing out that good out-of-sample forecast properties are an important but currently understudied criterion by which gravity equations should be evaluated.
Finally, we also contribute to the much smaller literature on the cross-sectional forecasting of international trade flows (see, for example, ITC, 2025). In contrast to this literature, our focus is on evaluating the out-of-sample predictive power of gravity equations, which we argue is much more important than its in-sample predictive power.
The rest of this paper is structured as follows. In Section 2, we demonstrate the conceptual link between the gravity equation’s out-of-sample forecasting performance and its reliability as a tool for estimating the causal impact of trade policy interventions. In Section 3, we introduce our general procedure for evaluating out-of-sample-predictions. In Section 4, we describe a number of statistical methods for forecasting trade flows, including the traditional and newer variants of the gravity equation, as well as our chosen machine learning methods. In Section 5, we use the proposed methodology to compare the out-of-sample forecasting performance of these methods. Finally, Section 6 concludes.
2 Gravity’s out-of-sample forecast performance and causal estimates of trade policy interventions
To illustrate the link between the out-of-sample forecast performance of gravity equations and their ability to estimate the causal effects of trade policy interventions, consider the following stripped-down model for trade flows ($y_{ijt}$) between two countries $i$ and $j$:
$E\left(y_{ijt}|D_{ijt}\right)=y_{ijt}^{0}e^{\tau D_{ijt}},$
where $D_{ijt}$ is the treatment variable we are interested in (e.g., an RTA dummy), and $y_{ijt}^{0}=E\left(y_{ijt}|D_{ijt}=0\right)$ is the expected value of trade in the absence of treatment. Accordingly, the expected value of trade in the presence of treatment is $E\left(y_{ijt}|D_{ijt}=1\right)=y_{ijt}^{0}e^{\tau}$, where $e^{\tau}$ is thus the causal effect of the treatment, which we want to estimate.
Within this general framework, an imputation estimator for $e^{\tau}$ can be constructed as (see Moreau-Kastler, 2025, and the references therein):
$e^{\hat{\tau}}=\frac{\sum_{D=1}y_{ijt}}{\sum_{D=1}\hat{y}_{ijt}^{0}},$ (1)
where the summation in both the numerator and denominator is across all observations with $D=1$, i.e., the “treated” observations. Furthermore, $\hat{y}_{ijt}^{0}$ is the predicted value of $y_{ijt}^{0}$ based on the model and estimator being used. If the sum in the denominator $\sum_{D=1}\hat{y}_{ijt}^{0}$ consistently estimates $\sum_{D=1}y_{ijt}^{0}$, expression (1) gives a consistent estimate for the multiplicative treatment effect $e^{\tau}$. Along the same lines, less biased predictions for each $y_{ijt}^{0}$ will result in less biased estimates of the treatment effect, and predictions with less variance will result in increased precision.
Note that for pairs that currently have an RTA in place (i.e., where $D_{ijt}=1$), $y_{ijt}^{0}$ is a counterfactual quantity that is not currently observed. This means that any evaluation of the suitability of a given estimator of $y_{ijt}^{0}$ should be based on its ability to predict trade flows out-of-sample. This is because an estimator might have good predictive performance in-sample, but standard issues of overfitting might imply that it will perform poorly for forecasting observations not contained in the underlying estimation sample, such as $y_{ijt}^{0}$.
The above explanation clarifies why good out-of-sample forecast performance is important for obtaining consistent estimates of the effects of trade policy interventions. While this is the main focus of our paper, researchers might of course also be interested in accurate out-of-sample forecasts for other reasons. For example, they might want to predict the trade potential between two countries not currently trading certain products with each other, as is the case for the ITC’s export potential assessment tool previously mentioned (see ITC, 2025). In both cases, the key question is the accuracy of the out-of-sample predictions for $y_{ijt}^{0}$ produced by different models. However, there is also a subtle difference between these two uses of out-of-sample predictions. As seen from the denominator of expression (1), if we are interested in obtaining consistent estimates of $e^{\tau}$, it is sufficient that the out-of-sample predictions for $y_{ijt}^{0}$ are correct on average, i.e., so over- and under-predictions can net out. By contrast, if we are interested in forecasting individual trade flows, such netting out of forecast errors is not possible. With this distinction in mind, we now describe our procedure for evaluating out-of-sample predictions.
3 Evaluating out-of-sample predictions
To evaluate the out-of-sample fit generated by different models for $y_{ijt}^{0}$, we proceed in five steps: [^6] The machine learning methods we use are not models in the usual sense but, to simplify the exposition, we often use the word ‘models’ to refer both to the gravity equations and the machine learning methods.
  1. 1.
    Select a model that we believe accurately represents the data generating process for $y_{ijt}^{0}$. For example, this could be the traditional gravity model, which postulates that trade flows between two countries $i$ and $j$ are a function of the two countries’ GDPs and the bilateral distance between them.
  2. 2.
    Split the relevant dataset into a training and a test sample; we describe how we split the sample in Section 5.3.
  3. 3.
    Use the training sample to estimate the chosen model.
  4. 4.
    Using the estimated model from step 3, compute predictions $\hat{y}_{ijt}^{0}$ for the observations in the test sample.
  5. 5.
    Compute measures of out-of-sample prediction performance for the estimators based on $K$ repetitions of steps 2 to 4; we discuss suitable measures of performance in Section 5.1.
4 Statistical models for trade flows
We compare the out-of-sample fit of several models for trade flows, each leading to different predictions $\hat{y}_{ijt}^{0}$ and hence different out-of-sample fits. The first model we consider is the traditional gravity model:
$y_{ijt}=\exp\left(\alpha_{0}+\beta_{1}\ln GDP_{it}+\beta_{2}\ln GDP_{jt}\text{+}\beta_{3}\ln dist_{ij}+\tau D_{ijt}\right)\varepsilon_{ijt},$ (2)
where $y_{ijt}$ are exports from country $i$ to country $j$ at time $t$, GDPit and GDPjt are the GDPs of the exporting and importing country, $dist_{ij}$ is the distance between the two countries, $D_{ijt}$ is again a dummy variable for the presence of treatment, and $\epsilon_{ijt}$ is the error term. [^7] In this section we present only stylized versions of the gravity equations. The models we estimate in Section 5 include additional regressors. This model implies that trade flows in the absence of treatment are given by:
$y_{ijt}^{0}=\exp\left(\alpha_{0}+\beta_{1}\ln GDP_{it}+\beta_{2}\ln GDP_{jt}\text{+}\beta_{3}\ln dist_{ij}\right)\varepsilon_{ijt}.$
The second model we consider is the so-called 2-way gravity model, which augments the traditional gravity model by adding exporter-time and importer-time fixed effects:
$y_{ijt}=\exp\left(\alpha_{it}+\gamma_{jt}+\beta_{3}\ln dist_{ij}+\tau D_{ijt}\right)\varepsilon_{ijt}.$ (3)
The motivation for including exporter-time and importer-time fixed effects ($\alpha_{it}$ and $\gamma_{jt}$) comes from the observation by Anderson and van Wincoop (2003) that gravity equations need to control for so-called multilateral resistance. Note that the previously included GDP terms are now absorbed by the fixed effects and so have to be dropped.
The third model we consider is the 3-way gravity model, which is an extension of the 2-way gravity model and augments it by further including bilateral fixed effects:
$y_{ijt}=\exp\left(\alpha_{it}+\gamma_{jt}+\eta_{ij}+\tau D_{ijt}\right)\varepsilon_{ijt}.$ (4)
The inclusion of bilateral fixed effects ($\eta_{ij}$) dates back to Baier and Bergstrand (2007) who argue that these fixed effects are crucial for mitigating endogeneity bias arising from omitted variables that influence both the probability of countries entering into RTAs as well as the amount of trade between them. Specifically, countries may strategically form trade agreements based on various factors, such as political alignment and shared economic interests, many of which are constant over time and specific to each country pair. These time-invariant unobserved factors may introduce endogeneity that poses challenges for estimation but can be controlled for through the inclusion of bilateral fixed effects. Note that it is now no longer possible to include the bilateral distance term as it would be collinear with the bilateral fixed effects.
In addition to the 2-way and 3-way gravity models, we also consider a version of the gravity model which we call the 1-way model. This version only includes bilateral fixed effects but omits exporter-time and importer-time fixed effects:
$y_{ijt}=\exp\left(\beta_{1}\ln GDP_{it}+\beta_{2}\ln GDP_{jt}\text{+}\eta_{ij}+\tau D_{ijt}\right)\varepsilon_{ijt}.$ (5)
The purpose of including the 1-way model is to evaluate the importance of controlling for the endogeneity of RTA formation, which both the 1-way and the 3-way models do, but the traditional and 2-way gravity models do not. Note that since the 1-way model excludes exporter-time and importer-time fixed effects, we can again include the GDP terms from the traditional gravity model.
All four versions of the gravity equation reflect conventional parameterizations for trade flows that have been widely used in the literature, and the 2-way and 3-way models are considered to be “theory-based” because of how they account for multilateral resistance (see Head and Mayer, 2014). While these models provide a natural starting point, we also consider nonparametric methods from the machine learning literature. Though these additional methods are not grounded in models of international trade, they have the advantage of being more flexible and could thus, in principle, yield better out-of-sample forecasts than the conventional gravity models. All the machine learning estimators were implemented using the user-written pystacked Stata command (Ahrens et al., 2023), which in turn relies on Python’s scikit-learn’s machine learning algorithms, see Pedregosa et al. (2011).
Applying machine learning estimators to our data requires choosing the set of predictors to be used. We consider two cases. We start by using just the covariates used by the traditional gravity equation (e.g., GDPs and distance). Additionally, we also include as regressors the three sets of fixed effects estimated from the 3-way gravity model. [^8] The alternative of including the full set of dummies representing the fixed effects is computationally infeasible. Including the fixed effects ensures that the machine learning methods have access to the same information as the 3-way specification, enabling a fairer comparison between the parametric and nonparametric approaches.
The first machine learning technique we consider is random forests. This technique consists of growing a large number of de-correlated decision trees and averaging over the predictions made by the individual trees (see Hastie et al., 2008, chapter 15, for details). In our results tables below, we refer to this approach as RF, or as RF-FE when the random forests are augmented by fixed effects.
The second machine learning technique we use is gradient boosting (GradBoost), which sequentially fits shallow trees, updating weights to minimize prediction error (see Friedman, 2001). We again present one version using the estimated fixed effects from the 3-way gravity model among the predictors (GB-FE) and one version without these fixed effects (GB).
Third, we estimate a neural network. Here, we again use the full set of covariates used in the traditional gravity model, either on their own (NN) or in addition to the estimated fixed effects from the 3-way gravity model (NN-FE). In contrast to the random forest and gradient boosting estimators, which are implemented using the default options in pystacked (Ahrens et al., 2023), for the neural networks estimators we use the Poisson loss to avoid negative fitted values and standardize all predictors with the StdScaler pipeline because neural networks are sensitive to the scale of the regressors (see LeCun et al., 2012).
The fourth machine learning technique we use is stacked generalization (Wolpert, 1992). This is an ensemble method that combines the forecasts obtained using the three machine learning methods described earlier, by using them as explanatory variables in a non-linear least squares regression without an intercept, in which the estimated coefficients are restricted to be non-negative and to sum to one (see Breiman, 1996, and Hastie et al., 2008). This, again, is the default option in pystacked (Ahrens et al., 2023). As usual, we present results using only the standard regressors, labeled Ens, and results in which the fixed effects are added to the set of regressors, labeled E-FE.
Finally, we estimate a “3-way-ML” gravity specification, which augments the 3-way gravity model with the logarithm of the fitted values from the ensemble without fixed effects. This specification allows us to assess whether the parametric 3-way model can be improved by incorporating flexible, nonlinear predictions from the machine learning methods. [^9] Note that the ensemble can predict zero trade flows, which makes it impossible to include the log of those forecast values. To address this, we impose that the 3-way-ML model also predict zero trade whenever the ensemble without fixed effects does so. We acknowledge that this modeling choice is ad hoc and may limit the effectiveness of integrating machine learning predictions into the gravity framework.
5 Comparing out-of-sample forecasts
In this section, we compare the out-of-sample forecast performance of the different models from Section 4. We start by presenting measures of predictive accuracy in Section 5.1 and introduce our data in Section 5.2. We then present our simulation design in Section 5.3 and finally present our results in Section 5.4.
5.1 Measures of predictive accuracy
We consider a range of measures of predictive accuracy that allow us to compare different aspects of the performance of the out-of-sample predictions obtained with the various estimators being considered.
We discuss in more detail below how we select the observations to be predicted, but for now simply note that there will be $n$ such observations. These $n$ observations correspond to our test sample in machine learning terminology, with the remaining NnN-n observations in our data serving as the estimation (or training) sample, where $N$ is the total number of observations in our dataset. We repeat the procedure of splitting the data into test and estimation sample and calculating our measures of predictive accuracy $K=1,000$ times. We then compute averages of our measures across the $K$ repetitions.
We start by considering measures of out-of-sample predictive accuracy based on the value of the imputation estimator, defined as the ratio of the mean observed outcome to the mean predicted value for the $n$ observations not used for estimation. That is, letting $\sum_{D_{k}=1}$ denote the sum over the set of $n$ observations in the test sample in repetition $k$ of the simulation procedure, the imputation estimator for the $k$-th repetition takes the form
$IE_{k}=\frac{\sum_{D_{k}=1}y_{ijt}}{\sum_{D_{k}=1}\hat{y}_{ijt}},$
where $\hat{y}_{ijt}$ represents the predicted value of $y_{ijt}$ estimated without using the $n_{k}$ observations excluded in repetition $k$.
The first measure of out-of-sample predictive accuracy that we consider is the mean of $IE_{k}$ across the $K$ repetitions. Because the observations being predicted are not different from the ones used for estimation (i.e., they were not “treated”), good out-of-sample predictive performance is characterized by a mean value of $IE_{k}$ close to $1$. Note that such a mean value can be achieved even if values of $IE_{k}$ for individual repetitions are different from unity, as over- and under-predictions can net out across the full $K$ repetitions. Therefore, to gauge the variation in the estimates for $e^{\tau}$ across samples, we also report the standard deviation of $IE_{k}$ across the $K$ simulation repetitions, denoted $SE_{IE}$, as well as the mean squared error of $IE$, computed as
$MSE_{IE}=\frac{1}{K}\sum_{k=1}^{K}(IE_{k}-1)^{2}.$
Apart from these measures based on the imputation estimator, we also consider the following measures of the ability of the model to predict individual observations.
  1. 1.
    Mean Absolute Error (MAE). This is the sum of the absolute errors for each of the predicted observations across all $K$ repetitions, divided by the total number of predicted observations across all repetitions:
    $\text{MAE}=\frac{\sum_{k=1}^{K}\sum_{D_{k}=1}\left|\hat{y}_{ijt}-y_{ijt}\right|}{\sum_{k=1}^{K}n_{k}}.$
  2. 2.
    Relative Mean Absolute Error (RMAE). This is obtained by dividing the MAE by the mean of the observed values in the prediction sample:
    $\text{RMAE}=\frac{\text{MAE}}{\frac{1}{\sum_{k=1}^{K}n_{k}}\sum_{k=1}^{K}\sum_{D_{k}=1}y_{ijt}}.$
    Note that the RMAE is a percentage, and therefore its magnitude is meaningful, making it easier to interpret than the MAE. In our results below, we therefore focus on the RMAE rather than the MAE.
  3. 3.
    Relative Root Mean Squared Error (RRMSE). This metric is defined as the root mean squared error (RMSE) divided by the mean of the observed values in the prediction sample:
     $\text{ RRMSE }=\frac{\text{RMSE}}{\frac{1}{\sum_{k=1}^{K}n_{k}}\sum_{k=1}^{K}\sum_{D_{k}=1}y_{ijt}},$
    with the RMSE given by:
    $\mathrm{RMSE}=\sqrt{\frac{\sum_{k=1}^{K}\sum_{D_{k}=1}\left(\hat{y}_{ijt}-y_{ijt}\right)^{2}}{\sum_{k=1}^{K}n_{k}}}.$
    As with the RMAE, we report the RRMSE rather than the RMSE below because it is a percentage and therefore its magnitude is meaningful and easy to interpret. [^10] Note that in contrast to the MAE, the RMSE downweights smaller errors and gives greater weight to larger errors.
  4. 4.
    Finally, we report the mean of the out-of-sample $R^{2}$, defined as the squared correlation between the observed and predicted values of trade over the $\sum_{k=1}^{K}n_{k}$ observations not used for estimation.
In the following, we distinguish between “estimation” and “prediction” measures of accuracy. The former include the mean of $IE$, $MSE_{IE}$, and $SE_{IE}$, while the latter include the RMAE, the RRMSE, and the out-of-sample $R^{2}$.
The purpose of this categorization is to distinguish between the use of out-of-sample predictions for estimation versus forecasting. As discussed earlier, if the aim is to obtain consistent estimates of policy intervention effects, it is sufficient for the model’s predictions to be correct on average across many observations. In such cases, over- and under-predictions may cancel out, so a mean imputation estimator close to one indicates a consistent estimate even if individual predictions are noisy. By contrast, when the goal is forecasting (i.e., focusing on the accuracy of predictions for individual observations), such averaging is not possible. In this case, accuracy measures that penalize the absolute or squared errors for individual predictions, such as the RMAE or the RRMSE, are more informative.
5.2 Data
We implement our tests of the out-of-sample performance of the estimators described in Section 4 on a standard and publicly available dataset of bilateral trade flows, that of Yotov (2025), which in turn is based on the United Nation’s Comtrade database. [^11] These data can be downloaded here: https://yotoyotov.com/Gravity_Undergrads.html The dataset includes aggregate trade and gravity variables for the $100$ largest exporters in the world, covering $98.9\%$ of world exports, $97.7\%$ of world imports, and $98.3\%$ of world GDP between 1990-2023.
5.3 Simulation design
We select our test and training samples from the full dataset described in Section 5.2 using the following procedure:
  1. (a)
    We first create a balanced panel by only including pairs with $30$ observations between 1994-2023. This leaves us with $261,540$ observations out of the original $320,920$.
  2. (b)
    Second, we estimate a 3-way gravity model using the ppmlhdfe command of Correia et al. (2020), which implements Gourieroux et al.’s (1984) PPML estimator, as recommended by Santos Silva and Tenreyro (2006). The specification includes time-varying indicators for whether both countries in a pair are members of the European Union, whether they share a customs union, whether they have a regional trade agreement (RTA), and whether the importer has imposed a trade sanction on the exporter; these regressors will also be used in all subsequent regressions. The model also includes exporter-time, importer-time and bilateral fixed effects, whose estimated values are saved.
  3. (c)
    We generate selection probabilities for each bilateral country pair by computing
    $p_{ij}=1/\left(1+\exp\left(a-b\hskip 0.85355pt\tilde{\eta}_{ij}\right)\right),$
    where $\tilde{\eta}_{ij}$ are the saved bilateral fixed effects estimates from the 3-way regression in (b), standardized to have mean zero and standard deviation one, $a$ is a parameter that controls the number of observations to be predicted, and $b$ controls the degree to which the probability of selection depends on time-invariant pair characteristics. We consider two samples, corresponding to different values of $a$ and $b$:
    1. (i)
      By selecting $a=5$ and $b=1$, our first sample is constructed such that a pair $ij$ is selected for prediction if $u_{ij}<p_{ij}$, where $u_{ij}\sim U(0,1)$ is drawn independently for each pair. For each selected pair, we then choose the most recent observations for prediction. Specifically, we select all years after the 6th decile of observed years for that pair, plus a discrete random variable uniformly distributed between 2-2 and $2$. We refer to this sample as the “endogenous treatment” case (see Table 1 below). It is endogenous in the sense that pairs with large positive $\tilde{\eta}_{ij}$ are more likely to be selected, reflecting the endogeneity commonly observed in the formation of RTAs, as discussed by Baier and Bergstrand (2007) and in Section 4.
    2. (ii)
      With $a=4.6$ and $b=0$, our second sample is of comparable size to the first, but now the probability of selection is independent of the bilateral fixed effects (the rest of the sample selection procedure is as before). We refer to this as the “exogenous treatment” sampling case (see Table 2 below). While we believe that case (i) is the most empirically relevant setting, assuming endogeneity naturally helps the forecast performance of models including bilateral fixed effects, such as the 3-way-gravity model. Thus, we also present results without endogeneity in order to see how the different models are affected by endogeneity.
Once the observations to be predicted are selected, we allocate the remaining observations from the sample from part (a) above to our estimation sample and estimate the statistical models outlined in Section 4.
Note that this potentially involves splitting the estimation sample into further estimation and test samples. This is because some of our machine learning methods use cross validation-based regularization procedures to choose their underlying model parameters, and this is done by minimizing the out-of-sample prediction error on a secondary test dataset. But we stress that the original test sample chosen in step c) above is not available for this tuning procedure, ensuring that our prediction exercise is purely out-of-sample. Of course, the gravity models we consider do not require tuning, so no further split of the original estimation sample is required for those.
5.4 Results
Table 1 shows results for the case of endogenous treatment (i.e., with $b=1$) which, across the $1,000$ replications, yields group sizes ($n_{k}$) ranging from $680$ to $1354$, with a mean of $1018.5$.
Table 1: Results for the case of endogenous treatment
Traditional 2-way 1-way 3-way Ens RF GB NN E-FE RF-FE GB-FE NN-FE 3-way-ML
Estimation
Mean $IE_{k}$} 1.028 1.154 0.964 0.993 1.007 1.007 1.069 1.039 1.031 1.042 1.007 1.005 1.037
$SE_{IE}$ 0.217 0.238 0.107 0.093 0.177 0.177 0.221 0.217 0.087 0.095 0.139 0.108 0.180
$MSE_{IE}$ 0.048 0.080 0.013 0.009 0.031 0.031 0.054 0.049 0.009 0.011 0.019 0.012 0.034
Prediction
RMAE 0.562 0.477 0.318 0.260 0.432 0.432 0.566 0.526 0.243 0.258 0.425 0.263 0.404
RRMSE 2.360 2.310 1.840 1.559 2.348 2.348 2.448 2.245 1.026 1.217 1.554 1.280 2.395
$R^{2}$ 0.718 0.781 0.890 0.914 0.791 0.791 0.718 0.747 0.913 0.906 0.817 0.913 0.807
Notes: This table shows simulation results for the case of endogenous treatment. The “Estimation” part focuses on the goal of obtaining consistent estimates of the effects of policy interventions (average forecasts), whereas the “Prediction” part focuses on the accuracy of forecasts for individual observations. The first four columns refer to gravity models (Traditional, 2-way, 1-way, 3-way). The remaining columns refer to machine learning methods. See the text for details.
Comparing the four gravity models first, we see that the 3-way model performs best by far. It yields a mean imputation estimator closest to one ($0.993$), has the smallest $SE_{IE}$ ($0.093$), and the lowest $MSE_{IE}$ ($0.009$). The 3-way model also achieves the lowest RMAE ($0.260$) and RRMSE ($1.559$), and the highest average out-of-sample $R^{2}$ ($0.914$).
Also of interest is the fact that the 2-way model performs worse than the 1-way and the traditional model. The most likely explanation for its poor performance relative to the 1-way model is that it does not control for the endogeneity of the treatment, because it does not include bilateral fixed effects. Interestingly, however, the traditional gravity model also sometimes outperforms the 2-way model even though it does not control for bilateral fixed effects either. A plausible explanation for the latter finding is that not controlling for the multilateral resistance terms that motivate the country-time fixed effects used in the 2-way and 3-way models induces a downward bias that partially offsets the upward bias from not accounting for the endogeneity of the treatment.
Moving on to the machine learning methods, we find that some methods can do as well, or even better, than the 3-way model in terms of the mean of $IE_{k}$ (Ens, RF, GB-FE, and NN-FE), but have much higher $SE_{IE}$ (see the estimation part of the table). In turn, E-FE (the ensemble method augmented with the estimated fixed effects from the 3-way model) has the smallest $SE_{IE}$ and a mean of $IE_{k}$ reasonably close to $1$. As a result, the 3-way and E-FE estimators have virtually identical values of the $MSE_{IE}$ ($0.009$) and dominate all other methods according to this criterion.
The one area where the machine learning methods can outperform the 3-way model is in the prediction of individual observations, which is the bottom part of the table. Here, the E-FE approach comes out ahead of the 3-way model both for the RMAE and RRMSE, although not for the out-of-sample $R^{2}$. However, the 3-way model outperforms almost all other machine learning methods in terms of the RMAE, as well as in terms of the out-of-sample $R^{2}$.
Finally, we note that the 3-way-ML estimator performs worse than the simpler 3-way gravity equation. Although this deserves further scrutiny, we conjecture that the deterioration in performance is due to overfitting. In any case, this is a reassuring result in the sense that it suggests that there is little reason to consider models more complex than the 3-way gravity equation.
We thus conclude that if a researcher is interested in obtaining consistent estimates of trade policy interventions such as RTAs, the 3-way gravity model is hard to beat. It is only when the researcher is interested in predicting individual trade flows out-of-sample that machine learning techniques can improve upon the performance of the 3-way model. However, this improvement is only possible if these techniques take account of the information from the fixed effects; the machine learning models not making use of these fixed estimates do not offer such an improvement. [^12] However, they do often outperform the traditional gravity equation that uses the same information set.
Table 2 shows results for the case with exogenous treatment (i.e., with $b=0$) which, across the $1,000$ replications, yields group sizes ($n_{k}$) ranging from $718$ to $1,399$, with a mean of $1044.9$.
Table 2: Results for the case of exogenous treatment
Traditional 2-way 1-way 3-way Ens RF GB NN E-FE RF-FE GB-FE NN-FE 3-way-ML
Estimation
Mean $IE_{k}$} 0.883 1.005 0.972 1.003 0.922 0.922 0.918 0.904 1.020 1.026 0.961 1.012 0.982
$SE_{IE}$ 0.234 0.242 0.140 0.115 0.210 0.210 0.251 0.236 0.113 0.120 0.180 0.133 0.212
$MSE_{IE}$ 0.068 0.059 0.020 0.013 0.050 0.050 0.070 0.065 0.013 0.015 0.034 0.018 0.045
Prediction
RMAE 0.639 0.519 0.330 0.268 0.491 0.491 0.670 0.602 0.255 0.270 0.477 0.274 0.440
RRMSE 3.348 3.299 2.565 2.158 3.435 3.435 3.480 3.221 1.460 1.711 2.152 1.756 3.526
$R^{2}$ 0.701 0.782 0.882 0.906 0.773 0.773 0.690 0.720 0.903 0.896 0.795 0.904 0.800
Notes: This table shows simulation results for the case of exogenous treatment. The “Estimation” part focuses on the goal of obtaining consistent estimates of the effects of policy interventions (average forecasts), whereas the “Prediction” part focuses on the accuracy of forecasts for individual observations. The first four columns refer to gravity models (Traditional, 2-way, 1-way, 3-way). The remaining columns refer to machine learning methods. See the text for details.
The results from Table 2 largely confirm those from Table 1.131313We also consider a case with $a=7.5$ and $b=1$, which leads to values of $n_{k}$ between $10$ and $203$, with a mean of $84.2$, and the results are qualitatively similar to those in Table 1. The 3-way model now outperforms all competitors in terms of the mean of the imputation estimator. However, the E-FE estimator beats it in terms of $SE_{IE}$, and the two estimators have again virtually indistinguishable values of $MSE_{IE}$. The one notable difference to Table 1 is the better performance of the 2-way model, which is now comparable to the 3-way model in terms of the bias of the imputation estimator and outperforms both the traditional gravity model and the 1-way model in this respect. This confirms our earlier supposition that the endogeneity of the treatment explains the poor performance of the 2-way model in Table 1. Turning now to the bottom part of the table, the E-FE approach is again the best performing method, both for the RMAE and RRMSE but not for the out-of-sample $R^{2}$; the 3-way model outperforms all other machine learning methods in terms of the RMAE and out-of-sample $R^{2}$.
6 Conclusion
In this paper, we argue that out-of-sample predictive performance of gravity equations is closely linked to their ability to yield consistent estimates of the causal effects of policy interventions such as RTAs. We present a methodological approach to compare the predictive performance of different estimation methods, and we use it to evaluate the reliability of estimates obtained with four different types of gravity equations (traditional, 1-way, 2-way and 3-way), as well as with atheoretical forecasting methods from the machine learning literature.
We find that the 3-way gravity model is difficult to beat when the purpose is to obtain consistent estimates of the effects of policy interventions. The only area where machine learning methods can outperform the 3-way model is the prediction of individual trade flows. Given that the evaluation of policy interventions is arguably the most important reason for the use of the gravity equation in applied international trade research, we see this result as reassuring, in the sense that it further cements the role of the 3-way gravity model as the state-of-the-art specification for gravity estimation. Future research will hopefully adopt our methodological approach and investigate whether our results hold in different datasets and using different machine learning estimators.

### v2

@@
 We propose a methodology that compares different versions of the gravity equation, both among themselves and with machine learning-based forecast methods such as random forests and neural networks.
-We find that the 3-way gravity model is difficult to beat in terms of out-of-sample average predictive performance, further justifying its place as the predominant tool for applied trade policy analysis.
+We find that the 3-way gravity model is difficult to beat in terms of out-of-sample average predictive performance, especially if a flexible specification is used.
+This result further justifies its place as the predominant tool for applied trade policy analysis.
 However, when the goal is to predict individual bilateral trade flows, the 3-way model can be outperformed by an ensemble machine learning method.
@@
 Gravity equations are by far the most popular tool in international trade for evaluating the effects of policy interventions such as regional trade agreements (RTAs).
-In its most basic form, the gravity model states that exports from country $i$ to country $j$ at time $t$ are a log-linear function of exporter and importer characteristics (e.g., GDP) and bilateral variables such as distance.
+In its most basic form, the gravity model states that exports from country $i$ to country $j$ at time $t$ are a log-linear function of exporter and importer characteristics (e.g., GDP) and bilateral variables (e.g., distance).
 The effects of RTAs and other policy interventions are usually evaluated by including an additional dummy variable for whether two countries $i$ and $j$ have an RTA in place.
@@
 This often requires making such forecasts, for example to predict the potential trade flows for a given pair of countries.
-[^5] For example, the International Trade Centre has developed a so-called ‘export potential assessment methodology’ in order to help countries identify promising products for export promotion activities (see ITC, 2025).
+For example, the International Trade Centre has developed a so-called ‘export potential assessment methodology’ in order to help countries identify promising products for export promotion activities (see ITC, 2025).
 The more accurate the forecast of potential trade flows between two countries, the better such export promotion activities can be targeted.
@@
 Given that the 3-way model is already considered to be the state-of-the-art version of the gravity equation (see Yotov et al., 2016), we see our results as a further justification for its use.
+However, our results also suggest that there may be (modest) gains from using a flexible specification of the way pair characteristics with time variation enter the model.
 Our research contributes to several strands in the international trade literature.
@@
 where the summation in both the numerator and denominator is across all observations with $D=1$, i.e., the “treated” observations.
-Furthermore, $\hat{y}_{ijt}^{0}$ is the predicted value of $y_{ijt}^{0}$ based on the model and estimator being used.
+Furthermore, $\hat{y}_{ijt}^{0}$ is the predicted value of $y_{ijt}^{0}$ obtained using only the observations with $D=0$ and the selected model and estimator.
+[^5] Notice that the standard PPML estimator of $e^{\tau}$ is similar to (1), except that the values in the denominator are estimated using the full sample.
 If the sum in the denominator $\sum_{D=1}\hat{y}_{ijt}^{0}$ consistently estimates $\sum_{D=1}y_{ijt}^{0}$, expression (1) gives a consistent estimate for the multiplicative treatment effect $e^{\tau}$.
@@
   2. 2.
-    Split the relevant dataset into a training and a test sample; we describe how we split the sample in Section 5.3.
+    Split the relevant dataset into a training and a test sample; we describe how we split the data in Section 5.3.
   3. 3.
@@
 We start by using just the covariates used by the traditional gravity equation (e.g., GDPs and distance).
-Additionally, we also include as regressors the three sets of fixed effects estimated from the 3-way gravity model.
+Additionally, we also include as regressors the three sets of fixed effects estimated with the 3-way gravity model for the corresponding training sample.
 [^8] The alternative of including the full set of dummies representing the fixed effects is computationally infeasible.
@@
 As usual, we present results using only the standard regressors, labeled Ens, and results in which the fixed effects are added to the set of regressors, labeled E-FE.
-Finally, we estimate a “3-way-ML” gravity specification, which augments the 3-way gravity model with the logarithm of the fitted values from the ensemble without fixed effects.
+Finally, we estimate a “3-way-ML” gravity specification, which augments the 3-way gravity model with the logarithm of the fitted values from the neural network estimator without fixed effects, which are necessarily positive.
 This specification allows us to assess whether the parametric 3-way model can be improved by incorporating flexible, nonlinear predictions from the machine learning methods.
-[^9] Note that the ensemble can predict zero trade flows, which makes it impossible to include the log of those forecast values.
-To address this, we impose that the 3-way-ML model also predict zero trade whenever the ensemble without fixed effects does so.
-We acknowledge that this modeling choice is ad hoc and may limit the effectiveness of integrating machine learning predictions into the gravity framework.
+[^9] Note that the predictions from the other machine learning methods are not necessarily positive, and therefore cannot be used to augment the 3-way model in the same way.
 5 Comparing out-of-sample forecasts
@@
 We repeat the procedure of splitting the data into test and estimation sample and calculating our measures of predictive accuracy $K=1,000$ times.
-We then compute averages of our measures across the $K$ repetitions.
+We then compute descriptive statistics of our measures across the $K$ repetitions.
 We start by considering measures of out-of-sample predictive accuracy based on the value of the imputation estimator, defined as the ratio of the mean observed outcome to the mean predicted value for the $n$ observations not used for estimation.
@@
 [^11] These data can be downloaded here: https://yotoyotov.com/Gravity_Undergrads.html The dataset includes aggregate trade and gravity variables for the $100$ largest exporters in the world, covering $98.9\%$ of world exports, $97.7\%$ of world imports, and $98.3\%$ of world GDP between 1990-2023.
+Besides data on trade, the database includes standard explanatory variables that include a measure of distance, GDPs for importer and exporter, and indicators for membership of the European Union, membership of a customs union, regional trade agreements, contiguity, common language, and colonial ties, as well as an indicator for whether the importer has imposed a trade sanction on the exporter.
+All these regressors are used in all subsequent regressions, although in come cases they drop out due to collinearity with the fixed effects.
 5.3 Simulation design
@@
 (2020), which implements Gourieroux et al.’s (1984) PPML estimator, as recommended by Santos Silva and Tenreyro (2006).
-The specification includes time-varying indicators for whether both countries in a pair are members of the European Union, whether they share a customs union, whether they have a regional trade agreement (RTA), and whether the importer has imposed a trade sanction on the exporter; these regressors will also be used in all subsequent regressions.
-The model also includes exporter-time, importer-time and bilateral fixed effects, whose estimated values are saved.
+The specification includes all the regressors described above, as well as exporter-time, importer-time and bilateral fixed effects, whose estimated values are saved.
   3. (c)
@@
 Table 1: Results for the case of endogenous treatment
-Traditional 2-way 1-way 3-way Ens RF GB NN E-FE RF-FE GB-FE NN-FE 3-way-ML
+Trad 2-way 1-way 3-way Ens RF GB NN E-FE RF-FE GB-FE NN-FE 3-way-ML
 Estimation
-Mean $IE_{k}$} 1.028 1.154 0.964 0.993 1.007 1.007 1.069 1.039 1.031 1.042 1.007 1.005 1.037
-$SE_{IE}$ 0.217 0.238 0.107 0.093 0.177 0.177 0.221 0.217 0.087 0.095 0.139 0.108 0.180
-$MSE_{IE}$ 0.048 0.080 0.013 0.009 0.031 0.031 0.054 0.049 0.009 0.011 0.019 0.012 0.034
+Mean $IE_{k}$} 1.028 1.154 0.964 0.993 1.007 1.007 1.069 1.039 1.031 1.042 1.007 1.005 1.000
+$SE_{IE}$ 0.217 0.238 0.107 0.093 0.177 0.177 0.221 0.217 0.087 0.095 0.139 0.108 0.084
+$MSE_{IE}\times 10$ 0.480 0.805 0.127 0.086 0.314 0.314 0.538 0.485 0.085 0.107 0.192 0.117 0.070
 Prediction
-RMAE 0.562 0.477 0.318 0.260 0.432 0.432 0.566 0.526 0.243 0.258 0.425 0.263 0.404
-RRMSE 2.360 2.310 1.840 1.559 2.348 2.348 2.448 2.245 1.026 1.217 1.554 1.280 2.395
-$R^{2}$ 0.718 0.781 0.890 0.914 0.791 0.791 0.718 0.747 0.913 0.906 0.817 0.913 0.807
+RMAE 0.562 0.477 0.318 0.260 0.432 0.432 0.566 0.526 0.243 0.258 0.425 0.263 0.244
+RRMSE 2.360 2.310 1.840 1.559 2.348 2.348 2.448 2.245 1.026 1.217 1.554 1.280 1.114
+$R^{2}$ 0.718 0.781 0.890 0.914 0.791 0.791 0.718 0.747 0.913 0.906 0.817 0.913 0.917
 Notes: This table shows simulation results for the case of endogenous treatment.
@@
 The first four columns refer to gravity models (Traditional, 2-way, 1-way, 3-way).
-The remaining columns refer to machine learning methods.
+The remaining columns refer to machine-learning based methods.
 See the text for details.
@@
 Moving on to the machine learning methods, we find that some methods can do as well, or even better, than the 3-way model in terms of the mean of $IE_{k}$ (Ens, RF, GB-FE, and NN-FE), but have much higher $SE_{IE}$ (see the estimation part of the table).
-In turn, E-FE (the ensemble method augmented with the estimated fixed effects from the 3-way model) has the smallest $SE_{IE}$ and a mean of $IE_{k}$ reasonably close to $1$.
-As a result, the 3-way and E-FE estimators have virtually identical values of the $MSE_{IE}$ ($0.009$) and dominate all other methods according to this criterion.
+In turn, E-FE (the ensemble method augmented with the estimated fixed effects from the 3-way model) has a smaller $SE_{IE}$ than the 3-way model and a mean of $IE_{k}$ reasonably close to $1$.
+As a result, the 3-way and E-FE estimators have virtually identical values of the $MSE_{IE}$ ($0.009$).
 The one area where the machine learning methods can outperform the 3-way model is in the prediction of individual observations, which is the bottom part of the table.
@@
 However, the 3-way model outperforms almost all other machine learning methods in terms of the RMAE, as well as in terms of the out-of-sample $R^{2}$.
-Finally, we note that the 3-way-ML estimator performs worse than the simpler 3-way gravity equation.
-Although this deserves further scrutiny, we conjecture that the deterioration in performance is due to overfitting.
-In any case, this is a reassuring result in the sense that it suggests that there is little reason to consider models more complex than the 3-way gravity equation.
-We thus conclude that if a researcher is interested in obtaining consistent estimates of trade policy interventions such as RTAs, the 3-way gravity model is hard to beat.
-It is only when the researcher is interested in predicting individual trade flows out-of-sample that machine learning techniques can improve upon the performance of the 3-way model.
-However, this improvement is only possible if these techniques take account of the information from the fixed effects; the machine learning models not making use of these fixed estimates do not offer such an improvement.
-[^12] However, they do often outperform the traditional gravity equation that uses the same information set.
+Finally, we find that the 3-way-ML estimator outperforms all other estimators we considered in terms of estimation criteria, and has the best out-of-sample $R^{2}$.
+Since the model includes 3-way fixed effects, the improvement over the basic 3-way model can only come from the augmented model that better captures the effect of pair characteristics with time variation, such as regional trade agreements and customs unions.
+This result is not surprising because it is well known that these can have very heterogeneous effects (see, e.g., Baier et al., 2019), and we model them with single dummies.
+Therefore, our results suggest that, when the aim is to evaluate policy interventions, it may be useful to employ a flexible specification of the 3-way model that more directly captures the effect of time-varying pair characteristics.
+This can be done parametrically (for example, by including interactions of the standard dummies with partners’ characteristics), or nonparametrically as we do in the 3-way-ML estimator.
+[^12] The advantage of the nonparametric way of controlling for these variables is that it avoids the potential overfitting from rich parametric specifications.
+In any case, our results suggest that the gains from using more complex models is relatively limited.
+We thus conclude that if a researcher is interested in obtaining consistent estimates of trade policy interventions such as RTAs, the 3-way gravity model is hard to beat, especially if it is flexibly specified.
+It is only when the researcher is interested in predicting individual trade flows out-of-sample that machine learning techniques can improve upon the performance of the 3-way models.
+However, this improvement is marginal and is only possible if these techniques take account of the information from the fixed effects; the machine learning models not making use of these fixed estimates do not offer such an improvement.
+[^13] However, they often outperform the traditional gravity equation that uses the same information set.
 Table 2 shows results for the case with exogenous treatment (i.e., with $b=0$) which, across the $1,000$ replications, yields group sizes ($n_{k}$) ranging from $718$ to $1,399$, with a mean of $1044.9$.
 Table 2: Results for the case of exogenous treatment
-Traditional 2-way 1-way 3-way Ens RF GB NN E-FE RF-FE GB-FE NN-FE 3-way-ML
+Trad 2-way 1-way 3-way Ens RF GB NN E-FE RF-FE GB-FE NN-FE 3-way-ML
 Estimation
-Mean $IE_{k}$} 0.883 1.005 0.972 1.003 0.922 0.922 0.918 0.904 1.020 1.026 0.961 1.012 0.982
-$SE_{IE}$ 0.234 0.242 0.140 0.115 0.210 0.210 0.251 0.236 0.113 0.120 0.180 0.133 0.212
-$MSE_{IE}$ 0.068 0.059 0.020 0.013 0.050 0.050 0.070 0.065 0.013 0.015 0.034 0.018 0.045
+Mean $IE_{k}$} 0.883 1.005 0.972 1.003 0.922 0.922 0.918 0.904 1.020 1.026 0.961 1.012 1.002
+$SE_{IE}$ 0.234 0.242 0.140 0.115 0.210 0.210 0.251 0.236 0.113 0.120 0.180 0.133 0.111
+$MSE_{IE}\times 10$ 0.068 0.059 0.020 0.013 0.050 0.050 0.070 0.065 0.013 0.015 0.034 0.018 0.012
 Prediction
-RMAE 0.639 0.519 0.330 0.268 0.491 0.491 0.670 0.602 0.255 0.270 0.477 0.274 0.440
-RRMSE 3.348 3.299 2.565 2.158 3.435 3.435 3.480 3.221 1.460 1.711 2.152 1.756 3.526
-$R^{2}$ 0.701 0.782 0.882 0.906 0.773 0.773 0.690 0.720 0.903 0.896 0.795 0.904 0.800
+RMAE 0.639 0.519 0.330 0.268 0.491 0.491 0.670 0.602 0.255 0.270 0.477 0.274 0.259
+RRMSE 3.348 3.299 2.565 2.158 3.435 3.435 3.480 3.221 1.460 1.711 2.152 1.756 1.742
+$R^{2}$ 0.701 0.782 0.882 0.906 0.773 0.773 0.690 0.720 0.903 0.896 0.795 0.904 0.908
 Notes: This table shows simulation results for the case of exogenous treatment.
@@
 The first four columns refer to gravity models (Traditional, 2-way, 1-way, 3-way).
-The remaining columns refer to machine learning methods.
+The remaining columns refer to machine-learning based methods.
 See the text for details.
 The results from Table 2 largely confirm those from Table 1.141414We also consider a case with $a=7.5$ and $b=1$, which leads to values of $n_{k}$ between $10$ and $203$, with a mean of $84.2$, and the results are qualitatively similar to those in Table 1.
-The 3-way model now outperforms all competitors in terms of the mean of the imputation estimator.
-However, the E-FE estimator beats it in terms of $SE_{IE}$, and the two estimators have again virtually indistinguishable values of $MSE_{IE}$.
 The one notable difference to Table 1 is the better performance of the 2-way model, which is now comparable to the 3-way model in terms of the bias of the imputation estimator and outperforms both the traditional gravity model and the 1-way model in this respect.
 This confirms our earlier supposition that the endogeneity of the treatment explains the poor performance of the 2-way model in Table 1.
-Turning now to the bottom part of the table, the E-FE approach is again the best performing method, both for the RMAE and RRMSE but not for the out-of-sample $R^{2}$; the 3-way model outperforms all other machine learning methods in terms of the RMAE and out-of-sample $R^{2}$.
 6 Conclusion
@@
 We find that the 3-way gravity model is difficult to beat when the purpose is to obtain consistent estimates of the effects of policy interventions.
+However, our results suggest that there may be moderate gains, especially in precision, from allowing pair characteristics with time variation to enter the model flexibly.
 The only area where machine learning methods can outperform the 3-way model is the prediction of individual trade flows.